how do you solve the problem ____, 108,_____,162,_____, 270,____,
Answer by stanbon(75887)

solve the problem ____, 108,_____,162,_____, 270,____,
----
Note:: 162-108 = 54
and 270-162 = 108 = 2*54
-----
1st: 81
2nd: 81+27 = 108
3rd: 108+27 = 135
4th: 135+27 = 162
5th: 162+27 = 189
6th: 189+27 = 216
7th: 216+27 = 243
8th: 243+27 = 270
==========
Cheers,
Stan H.
-----------